# SAM 2: Segment Anything in Images and Videos

### v1

@@ D.2.1 Annotation protocol
D.2.2 Data engine phase comparison
@@
Appendix E Details on zero-shot transfer experiments
@@
E.1 Zero-shot video tasks
E.1.1 Video dataset details
@@
Among these 17 zero-shot video datasets above, 9 of them (EndoVis, ESD, LVOSv2, LV-VIS, UVO, VOST, PUMaVOS, Virtual KITTI 2, and VIPSeg) have dense object segments annotated for every video frame. In the remaining 8 datasets (Wildfires, VISOR, FBMS, Ego-Exo4D, Cityscapes, Lindenthal Camera, HT1080WT Cells, and Drosophila Heart), the object segments are sparsely annotated over only a subset of video frames, and we compute the metrics on those frames where the ground-truth segmentation masks are available. In most evaluations of the paper, we only evaluate zero-shot performance on the 9 densely annotated datasets, while in our semi-supervised VOS evaluation (§6.1.2), we evaluate on all these 17 datasets listed above.
E.1.2 Interactive offline and online evaluation details
@@
In both settings, we evaluate on 9 densely annotated datasets in §E.1.1 (EndoVis, ESD, LVOSv2, LV-VIS, UVO, VOST, PUMaVOS, Virtual KITTI 2, and VIPSeg). If a video contains multiple objects to segment in its ground-truth annotations, we perform inference on each object independently. We simulate interactive video segmentation with $N_{\mathrm{click}}=3$ clicks per frame, assuming that the user would visually locate the object to label it (with initial clicks) or to refine the current segmentation prediction of it (with correction clicks). Specifically, when starting the first pass (where there are not any existing predictions yet), we place an initial click on the first frame at the center [^1] The center of a mask is defined as the mask pixel that has the largest Euclidean distance to the mask boundary. of the object’s ground-truth mask and then interactively add two more clicks based on the center of the error region (between the ground-truth mask and the predicted segments on the first frame). Then in subsequent passes (where there are already predicted segments), we interactively add three clicks based on the center of the error region (between the ground-truth mask and the predicted segments on the frame being prompted).
@@
E.1.3 Semi-supervised VOS evaluation details
@@
Figure 12: Zero-shot performance of SAM 2 vs baselines (SAM+XMem++ and SAM+Cutie) under interactive offline evaluation with different numbers of interacted frames, using 3 clicks per interacted frame. See §E.1.2 for details.
@@
Figure 13: Zero-shot performance of SAM 2 vs baselines (SAM+XMem++ and SAM+Cutie) under interactive online evaluation with different numbers of interacted frames, using 3 clicks per interacted frame. See §E.1.2 for details.
In §6.1.2, we also compare with previous video tracking methods under the semi-supervised VOS setting (Pont-Tuset et al., 2017), where prompts (which can be foreground/background clicks, bounding boxes, or ground-truth object masks) are provided only on the first frame of the video. When using click prompts, we interactively sample either 1, 3 or 5 clicks on the first video frame, and then track the object based on these clicks. Following the click-based evaluation in prior work (Kirillov et al., 2023; Sofiiuk et al., 2022), the initial click is placed on the object center and subsequent clicks are obtained from the center of the error region.
@@
In this setting, we evaluate on all 17 zero-shot video datasets in §E.1.1. If a dataset does not follow the standard VOS format, we preprocess it into a format similar to MOSE (Ding et al., 2023). During processing, we ensure that all objects in each video have a valid non-empty segmentation mask on the first frame to be compatible with semi-supervised VOS evaluation. In case an object doesn’t appear in the first frame, we create a separate video for it starting from the first frame where the object appears.
@@
Figure 14: Zero-shot performance on 17 video datasets of SAM 2 vs two baselines (SAM+XMem++ and SAM+Cutie) under semi-supervised VOS evaluation using different prompts (1, 3 or 5 clicks, bounding boxes, or ground-truth masks on the first video frame), with the averaged performance across datasets for each type of prompt shown in Table 4 in the main text. See §E.1.3 for details.
E.1.4 SAM+XMem++ and SAM+Cutie baseline details
@@
E.2 DAVIS interactive benchmark
We also evaluate on the DAVIS interactive benchmark (Caelles et al., 2018), which resembles our interactive offline evaluation in §6.1.1, where in each round of interaction, the evaluation server would provide new annotations on frames with the worst segmentation performance. The official DAVIS eval toolkit provides scribble prompts during interactions, while other work such as CiVOS (Vujasinović et al., 2022) has also extended this to cover click prompts.
Here we follow CiVOS to use positive and negative clicks as input prompts and adopt the same strategy for click sampling. We report the $\mathcal{J}\&\mathcal{F}$@60s and AUC-$\mathcal{J}\&\mathcal{F}$ metrics on this benchmark as provided by its evaluator, and compare to two baselines: MiVOS (Cheng et al., 2021b), which directly uses the provided scribbles via a scribble-to-mask module (and is also extended to click prompts in Vujasinović et al. (2022)), and CiVOS, which samples click from the provided scribbles. The results are shown in Table 14, where SAM 2 (based on click inputs) outperforms both baselines under click inputs. We note that SAM 2 often tends to segment object parts (e.g. a person’s arm) on the first click while the DAVIS dataset mainly contains whole objects (e.g. an entire person), which could penalize SAM 2’s $\mathcal{J}$$\mathcal{F}$ performance on this benchmark. We verified this by observing better accuracy (0.86 AUC-$\mathcal{J}$$\mathcal{F}$ and 0.89 $\mathcal{J}$$\mathcal{F}$@60s with click input) for an earlier model trained on fewer part annotations.
@@ E.2 DAVIS interactive benchmark
SAM 2 clicks 0.84 0.87
Table 14: Performance of SAM 2 and other models on the DAVIS interactive benchmark. For SAM 2, we use clicks as inputs following the click sampling strategy from CiVOS (Vujasinović et al., 2022). See §E.2 for details (: performance reported in Vujasinović et al. (2022)).
E.3 Zero-shot image tasks
E.3.1 Dataset details
For the interactive segmentation task, we evaluated SAM 2 on a comprehensive suite of 37 datasets. This suite includes the 23 datasets previously used by SAM for zero-shot evaluation. For completeness, we list the 23 datasets: LVIS (Gupta et al., 2019), ADE20K (Zhou et al., 2019), Hypersim (Roberts et al., 2021), Cityscapes (Cordts et al., 2016), BBBC038v1 (Caicedo et al., 2019), DOORS (Pugliatti & Topputo, 2022), DRAM (Cohen et al., 2022), EgoHOS (Zhang et al., 2022), GTEA (Fathi et al., 2011; Li et al., 2015), iShape (Yang et al., 2021a), NDD20 (Trotter et al., 2020), NDISPark (Ciampi et al., 2021, 2022), OVIS (Qi et al., 2022), PPDLS (Minervini et al., 2016), Plittersdorf (Haucke et al., 2022), STREETS (Snyder & Do, 2019), TimberSeg (Fortin et al., 2022), TrashCan (Hong et al., 2020), VISOR (Darkhalil et al., 2022; Damen et al., 2022), WoodScape (Yogamani et al., 2019), PIDRay (Wang et al., 2021a), ZeroWaste-f (Bashkirova et al., 2022), and IBD (Chen et al., 2022). For more detailed information about these datasets, we refer the reader to Kirillov et al. (2023). In addition to these 23 datasets, we evaluated on frames sampled from 14 video datasets to assess SAM 2’s performance on images from the video domain. The video datasets used are listed as follows: Lindenthal Camera Traps (LCT) (Haucke & Steinhage, 2021), VOST (Tokmakov et al., 2022), LV-VIS (Wang et al., 2023), FBMS (Brox et al., 2010), Virtual KITTI 2 (Cabon et al., 2020), Corsican Fire Database (CFD) (Toulouse et al., 2017), VIPSeg (Miao et al., 2022), Drosophila Heart OCM (DH OCM) (Fishman et al., 2023), EndoVis 2018 (Allan et al., 2020), ESD (Huang et al., 2023), UVO (Wang et al., 2021b), Ego-Exo4d (Grauman et al., 2023), LVOSv2 (Hong et al., 2024), and HT1080WT (Gómez-de Mariscal et al., 2021). Table 16 has a more detailed description of these datasets. (Some of these datasets are obtained from the same data source as the zero-shot video datasets in §E.1.1.)
E.3.2 Detailed zero-shot experiments
@@
Appendix F Details on comparison to state-of-the-art in semi-supervised VOS
@@
SAM 2 (Hiera-B+) 73.6 70.4 76.9 74.1 70.6 77.5
SAM 2 (Hiera-L) 75.6 72.3 78.9 77.6 74.0 81.1
SAM 2 (Hiera-T) 73.7 70.3 77.1 75.0 71.5 78.5
SAM 2 (Hiera-S) 72.7 69.4 76.0 74.9 71.4 78.4
SAM 2 (Hiera-B+) 75.3 71.8 78.7 74.7 71.2 78.2
SAM 2 (Hiera-L) 76.1 72.9 79.2 76.0 72.6 79.3
@@
SAM 2 (Hiera-B+) 74.9 70.2 79.6
SAM 2 (Hiera-L) 76.1 71.6 80.6
SAM 2 (Hiera-T) 73.6 68.8 78.3
SAM 2 (Hiera-S) 73.5 68.6 78.4
SAM 2 (Hiera-B+) 76.2 71.6 80.7
SAM 2 (Hiera-L) 77.9 73.1 82.7
@@
SwinB-DeAOT (Yang & Yang, 2022) 63.9 61.5 69.0 58.4 66.6
@@
SAM 2 (Hiera-B+) 75.8 78.9 85.3 65.1 73.8
SAM 2 (Hiera-L) 78.1 78.9 85.2 69.7 78.6
SAM 2 (Hiera-T) 75.3 75.2 81.8 67.7 76.4
SAM 2 (Hiera-S) 76.4 77.1 84.0 67.5 77.0
SAM 2 (Hiera-B+) 75.8 77.0 83.4 67.0 75.6
SAM 2 (Hiera-L) 79.8 80.0 86.6 71.6 81.1
@@
SAM 2 (Hiera-B+) 75.8 71.8 79.9 90.9 87.7 94.1 88.3 85.0 91.5 88.4 87.1 91.7 83.3 91.4
SAM 2 (Hiera-L) 77.2 73.3 81.2 91.6 88.3 94.9 89.0 85.8 92.2 89.1 87.5 92.0 84.5 92.4
SAM 2 (Hiera-T) 70.9 66.7 75.2 89.2 85.7 92.7 86.7 83.3 90.0 87.4 85.5 90.0 83.0 91.2
SAM 2 (Hiera-S) 71.5 67.3 75.6 88.8 85.4 92.2 86.3 82.8 89.9 87.5 85.7 90.1 83.2 91.2
SAM 2 (Hiera-B+) 72.8 68.8 76.9 88.9 85.5 92.2 86.7 83.2 90.1 87.9 86.2 90.7 83.2 91.4
SAM 2 (Hiera-L) 74.6 70.6 78.6 89.3 85.7 92.8 88.1 84.6 91.6 88.6 86.6 91.3 84.2 92.2
@@
Appendix G Model, data and annotation cards
G.1 Model card
@@
Repository https://github.com/facebookresearch/segment-anything-2
@@
Caveats and recommendations See Appendix B for limitations.
@@
Data sources See Appendix E
@@
Data source SAM 2 was trained on the SA-V dataset alongside internally available licensed video data. See Section 5 of the main text for more details and Appendix G.2 for the SA-V dataset data card.
@@
Data See Section 5 for more details about the SAM 2 training data. In Section 3 we show a geographic distribution of the videos and demographic distribution of the crowdworkers who collected the videos in the SA-V dataset.
@@
Risks and harms In Section 6.1.3 of the main text we analyze SAM 2 performance on people across demographic groups. When using SAM 2 in new settings, we suggest that researchers perform their own fairness evaluation for SAM 2 specific to their use case.
@@
G.2 Dataset card for SA-V dataset

### v2

@@ E.2.1 Annotation protocol
E.2.2 Data engine phase comparison
@@
Appendix F Details on zero-shot transfer experiments
@@
F.1 Zero-shot video tasks
F.1.1 Video dataset details
@@
Among these 17 zero-shot video datasets above, 9 of them (EndoVis, ESD, LVOSv2, LV-VIS, UVO, VOST, PUMaVOS, Virtual KITTI 2, and VIPSeg) have dense object segments annotated for every video frame. In the remaining 8 datasets (Wildfires, VISOR, FBMS, Ego-Exo4D, Cityscapes, Lindenthal Camera, HT1080WT Cells, and Drosophila Heart), the object segments are sparsely annotated over only a subset of video frames, and we compute the metrics on those frames where the ground-truth segmentation masks are available. In most evaluations of the paper, we only evaluate zero-shot performance on the 9 densely annotated datasets, while in our semi-supervised VOS evaluation (§6.2), we evaluate on all these 17 datasets listed above.
F.1.2 Interactive offline and online evaluation details
@@
In both settings, we evaluate on 9 densely annotated datasets in §F.1.1 (EndoVis, ESD, LVOSv2, LV-VIS, UVO, VOST, PUMaVOS, Virtual KITTI 2, and VIPSeg). If a video contains multiple objects to segment in its ground-truth annotations, we perform inference on each object independently. We simulate interactive video segmentation with $N_{\mathrm{click}}=3$ clicks per frame, assuming that the user would visually locate the object to label it (with initial clicks) or to refine the current segmentation prediction of it (with correction clicks). Specifically, when starting the first pass (where there are not any existing predictions yet), we place an initial click on the first frame at the center [^2] The center of a mask is defined as the mask pixel that has the largest Euclidean distance to the mask boundary. of the object’s ground-truth mask and then interactively add two more clicks based on the center of the error region (between the ground-truth mask and the predicted segments on the first frame). Then in subsequent passes (where there are already predicted segments), we interactively add three clicks based on the center of the error region (between the ground-truth mask and the predicted segments on the frame being prompted).
@@
F.1.3 Semi-supervised VOS evaluation details
@@
Figure 12: Zero-shot performance of SAM 2 vs baselines (SAM+XMem++ and SAM+Cutie) under interactive offline evaluation with different numbers of interacted frames, using 3 clicks per interacted frame. See §F.1.2 for details.
@@
Figure 13: Zero-shot performance of SAM 2 vs baselines (SAM+XMem++ and SAM+Cutie) under interactive online evaluation with different numbers of interacted frames, using 3 clicks per interacted frame. See §F.1.2 for details.
In §6.2, we also compare with previous video tracking methods under the semi-supervised VOS setting (Pont-Tuset et al., 2017), where prompts (which can be foreground/background clicks, bounding boxes, or ground-truth object masks) are provided only on the first frame of the video. When using click prompts, we interactively sample either 1, 3 or 5 clicks on the first video frame, and then track the object based on these clicks. Following the click-based evaluation in prior work (Kirillov et al., 2023; Sofiiuk et al., 2022), the initial click is placed on the object center and subsequent clicks are obtained from the center of the error region.
@@
In this setting, we evaluate on all 17 zero-shot video datasets in §F.1.1. If a dataset does not follow the standard VOS format, we preprocess it into a format similar to MOSE (Ding et al., 2023). During processing, we ensure that all objects in each video have a valid non-empty segmentation mask on the first frame to be compatible with semi-supervised VOS evaluation. In case an object doesn’t appear in the first frame, we create a separate video for it starting from the first frame where the object appears.
@@
Figure 14: Zero-shot performance on 17 video datasets of SAM 2 vs two baselines (SAM+XMem++ and SAM+Cutie) under semi-supervised VOS evaluation using different prompts (1, 3 or 5 clicks, bounding boxes, or ground-truth masks on the first video frame), with the averaged performance across datasets for each type of prompt shown in Table 4 in the main text. See §F.1.3 for details.
F.1.4 SAM+XMem++ and SAM+Cutie baseline details
@@
F.2 DAVIS interactive benchmark
We also evaluate on the DAVIS interactive benchmark (Caelles et al., 2018), which resembles our interactive offline evaluation in §6.1, where in each round of interaction, the evaluation server would provide new annotations on frames with the worst segmentation performance. The official DAVIS eval toolkit provides scribble prompts during interactions, while other work such as CiVOS (Vujasinović et al., 2022) has also extended this to cover click prompts.
Here we follow CiVOS to use positive and negative clicks as input prompts and adopt the same strategy for click sampling. We report the $\mathcal{J}\&\mathcal{F}$@60s and AUC-$\mathcal{J}\&\mathcal{F}$ metrics on this benchmark as provided by its evaluator, and compare to two baselines: MiVOS (Cheng et al., 2021b), which directly uses the provided scribbles via a scribble-to-mask module (and is also extended to click prompts in Vujasinović et al. (2022)), and CiVOS, which samples click from the provided scribbles. The results are shown in Table 14, where SAM 2 (based on click inputs) outperforms both baselines under click inputs. We note that SAM 2 often tends to segment object parts (e.g. a person’s arm) on the first click while the DAVIS dataset mainly contains whole objects (e.g. an entire person), which could penalize SAM 2’s $\mathcal{J}$$\mathcal{F}$ performance on this benchmark.
@@ F.2 DAVIS interactive benchmark
SAM 2 clicks 0.86 0.90
Table 14: Performance of SAM 2 and other models on the DAVIS interactive benchmark. For SAM 2, we use clicks as inputs following the click sampling strategy from CiVOS (Vujasinović et al., 2022). See §F.2 for details (: performance reported in Vujasinović et al. (2022)).
F.3 Zero-shot image tasks
F.3.1 Dataset details
For the interactive segmentation task, we evaluated SAM 2 on a comprehensive suite of 37 datasets. This suite includes the 23 datasets previously used by SAM for zero-shot evaluation. For completeness, we list the 23 datasets: LVIS (Gupta et al., 2019), ADE20K (Zhou et al., 2019), Hypersim (Roberts et al., 2021), Cityscapes (Cordts et al., 2016), BBBC038v1 (Caicedo et al., 2019), DOORS (Pugliatti & Topputo, 2022), DRAM (Cohen et al., 2022), EgoHOS (Zhang et al., 2022), GTEA (Fathi et al., 2011; Li et al., 2015), iShape (Yang et al., 2021a), NDD20 (Trotter et al., 2020), NDISPark (Ciampi et al., 2021, 2022), OVIS (Qi et al., 2022), PPDLS (Minervini et al., 2016), Plittersdorf (Haucke et al., 2022), STREETS (Snyder & Do, 2019), TimberSeg (Fortin et al., 2022), TrashCan (Hong et al., 2020), VISOR (Darkhalil et al., 2022; Damen et al., 2022), WoodScape (Yogamani et al., 2019), PIDRay (Wang et al., 2021a), ZeroWaste-f (Bashkirova et al., 2022), and IBD (Chen et al., 2022). For more detailed information about these datasets, we refer the reader to Kirillov et al. (2023). In addition to these 23 datasets, we evaluated on frames sampled from 14 video datasets to assess SAM 2’s performance on images from the video domain. The video datasets used are listed as follows: Lindenthal Camera Traps (LCT) (Haucke & Steinhage, 2021), VOST (Tokmakov et al., 2022), LV-VIS (Wang et al., 2023), FBMS (Brox et al., 2010), Virtual KITTI 2 (Cabon et al., 2020), Corsican Fire Database (CFD) (Toulouse et al., 2017), VIPSeg (Miao et al., 2022), Drosophila Heart OCM (DH OCM) (Fishman et al., 2023), EndoVis 2018 (Allan et al., 2020), ESD (Huang et al., 2023), UVO (Wang et al., 2021b), Ego-Exo4d (Grauman et al., 2023), LVOSv2 (Hong et al., 2024), and HT1080WT (Gómez-de Mariscal et al., 2021). Table 16 has a more detailed description of these datasets. (Some of these datasets are obtained from the same data source as the zero-shot video datasets in §F.1.1.)
F.3.2 Detailed zero-shot experiments
@@
Appendix G Details on comparison to state-of-the-art in semi-supervised VOS
@@
SAM 2 (Hiera-B+) 76.8 73.4 80.1 77.0 73.5 80.5
SAM 2 (Hiera-L) 77.9 74.5 81.3 78.4 74.9 82.0
SAM 2 (Hiera-T) 75.2 71.7 78.7 76.5 72.9 80.1
SAM 2 (Hiera-S) 77.0 73.6 80.5 76.6 73.0 80.2
SAM 2 (Hiera-B+) 77.5 74.1 80.9 78.2 74.8 81.7
SAM 2 (Hiera-L) 78.6 75.3 82.0 79.5 76.0 83.0
@@
SAM 2 (Hiera-B+) 78.0 73.2 82.7
SAM 2 (Hiera-L) 78.0 73.2 82.7
SAM 2 (Hiera-T) 77.5 73.0 82.1
SAM 2 (Hiera-S) 77.3 72.3 82.2
SAM 2 (Hiera-B+) 77.7 73.1 82.4
SAM 2 (Hiera-L) 80.1 75.4 84.9
@@
SwinB-DeAOT-L (Yang & Yang, 2022) 63.9 61.5 69.0 58.4 66.6
@@
SAM 2 (Hiera-B+) 78.7 80.6 87.2 69.0 77.8
SAM 2 (Hiera-L) 79.6 81.0 87.4 70.4 79.8
SAM 2 (Hiera-T) 77.3 79.9 86.9 67.1 75.4
SAM 2 (Hiera-S) 78.3 79.2 85.9 69.0 79.0
SAM 2 (Hiera-B+) 78.2 80.5 87.2 68.1 76.9
SAM 2 (Hiera-L) 80.6 81.7 88.2 71.4 81.0
@@
SAM 2 (Hiera-B+) 76.6 72.6 80.6 90.2 87.0 93.4 87.9 84.7 91.1 88.6 87.1 91.6 83.9 91.9
SAM 2 (Hiera-L) 77.9 73.9 81.9 90.7 87.5 94.0 87.7 84.6 90.9 89.3 87.5 92.0 84.8 92.8
SAM 2 (Hiera-T) 71.8 67.4 76.1 89.4 85.8 92.9 86.9 83.4 90.3 87.4 85.7 90.1 82.7 91.1
SAM 2 (Hiera-S) 73.5 69.2 77.7 89.6 86.3 92.9 87.6 84.1 91.1 88.0 85.9 90.2 83.9 91.9
SAM 2 (Hiera-B+) 73.8 69.7 77.9 90.0 86.8 93.1 86.6 83.3 89.8 88.2 86.1 90.6 84.0 92.1
SAM 2 (Hiera-L) 74.6 70.6 78.6 90.2 86.9 93.4 88.9 85.3 92.5 88.8 86.5 91.0 84.7 92.8
@@
Appendix H Model, data and annotation cards
H.1 Model card
@@
Repository https://github.com/facebookresearch/sam2
@@
Caveats and recommendations See Appendix C for limitations.
@@
Data sources See Appendix F
@@
Data source SAM 2 was trained on the SA-V dataset alongside internally available licensed video data. See Section 5 of the main text for more details and Appendix H.2 for the SA-V dataset data card.
@@
Data See Section 5 for more details about the SAM 2 training data. In Section E.1 we show a geographic distribution of the videos and demographic distribution of the crowdworkers who collected the videos in the SA-V dataset.
@@
Risks and harms In Section E.1.1 of the main text we analyze SAM 2 performance on people across demographic groups. When using SAM 2 in new settings, we suggest that researchers perform their own fairness evaluation for SAM 2 specific to their use case.
@@
H.2 Dataset card for SA-V dataset